# Localization of Hybridized Surface Plasmon Modes on Random Gold Nanoparticle Assemblies

### v1

@@ I Introduction
Strongly disordered micro- and nanostructures of (noble) metals, such as random assemblies of gold nanoparticles (NPs) or thin random gold networks, are currently used/investigated [1] for various applications such as electromechanical gas and vapor sensors [2, 3] or flexible electrodes, e.g., for neural implants [4] or monitoring of blood vessels [5]. The latter exploits the presence/absence of strongly localized surface plasmon (LSP) resonances in a particular frequency band depending on the composition and geometry of the nanostructure. The localization stays in contrast to the formation of delocalized modes in periodic structures (lattice resonances, quasi-continuous plasmon bands), which give rise to prominent quality factors, strong dispersion, and ballistic transport, amongst others [6, 7, 8, 9]. Both phenomena originate from the coupling and hybridization of LSP resonances via electromagnetic interaction [7]. Localization of surface plasmons in strongly coupling random networks and NP assemblies has been intensely studied by various groups [10, 11, 12, 13], establishing the fundamental role of the disorder in the observed spatial localization and spectral distribution. Recently, we have demonstrated the localization of surface plasmons in ultrathin random gold networks, which increased significantly in an energy interval from 0.3 to 1.6 eV, and the absence of hybridized modes beyond [10]. In order to model the intriguing localization behavior, we described the random network as a random assembly of strongly oblate Au NPs (i.e., the inverted system in the sense that holes of the network correspond to NPs) by exploiting Babinet’s principle [14, 15, 16], which was then amenable to numerical solutions. Notwithstanding obtaining good agreement between experiment and simulation, the question of how the thickness and other geometric parameters affected the localization behavior remained open.

### v2

# Geometry Dependent Localization of Surface Plasmons on Random Gold Nanoparticle Assemblies

@@ I Introduction
Strongly disordered micro- and nanostructures of (noble) metals, such as random assemblies of gold nanoparticles (NPs) or thin random gold networks, are currently used for various applications such as electromechanical gas and vapor sensors [1, 2] or flexible electrodes, e.g., for neural implants [3] or monitoring of blood vessels [4]. The latter exploits the presence of strongly localized surface plasmon (LSP) resonances in a particular frequency band depending on the composition and geometry of the nanostructure. Localization contrasts with the formation of delocalized modes in periodic structures (such as lattice resonances and quasicontinuous plasmon bands), which exhibit high quality factors, strong dispersion, and ballistic transport, among other characteristics [5, 6, 7, 8]. Both phenomena originate from the coupling and hybridization of LSP resonances via electromagnetic interaction [6]. Localization of surface plasmons in strongly coupling random networks and NP assemblies has been intensely studied by various groups [9, 10, 11, 12], establishing the fundamental role of the disorder in the observed spatial localization and spectral distribution. Recently, we have demonstrated the localization of surface plasmons in ultrathin random gold networks, which increased significantly in an energy interval from 0.3 to 1.6 eV, and the absence of hybridized modes beyond [9]. In order to model the intriguing localization behavior, we described the random network as a random assembly of strongly oblate Au NPs (i.e., the inverted system in the sense that holes of the network correspond to NPs) by exploiting Babinet’s principle [13, 14, 15], which was then amenable to numerical solutions. Notwithstanding obtaining good agreement between experiment and simulation, the question of how the thickness and other geometric parameters affected the localization behavior remained open.